find the (3) use cylindrical Coordinates to Volume of the Solid bounded by Z=9-2²-9² and z=5. (3)
Answer #1

Let dar caso, y = 8 Sina and Z=2.. alin. (dx/dy) = r dro So we betureen have to find the volume z=9-82 and z=5 Here, r will b

> Ves (inte de da > v= g (2 e) do v (81 - 18) do 7 V - I do JV = ax culte units uolume is Hence, required GR inbic unite